# EmbAdvisor: Adaptive Cache Management for Sustainable LLM Serving

### v1

@@ 3. Performance and Carbon of LLM Caching
In this section, we analyze the performance and carbon emissions of caching in LLM serving, and summarize the observations as takeaways. We illustrate the results using the Llama 3 70B model (Meta, 2024) and the ShareGPT dataset (Hugging Face, 2023), evaluated on the server described in Table 1. We use LMCache (LMCache Team, 2025b) as the caching system, which saves context on 16 TB SSDs. The cache has been initialized with 200k prompts, and we record the performance and carbon emission of 500 prompts. Details of our evaluation methodology are provided in Section 6.1.
3.1. LLM Caching Performance
We first study the performance of context caching in LLM serving, by focusing on two factors that impact the benefits: context length and request rate.
@@ 3.1. LLM Caching Performance
Figure 5. (a) Average latency of prefill and decode under different req rates, and the speedup from caching. (b) Average percentage of prefill and decode latency in a prompt.
@@ 3.1.1. Impact of Context Length
Caching reduces the computation of existing context. Therefore, we first evaluate its benefits under different context lengths. Figure 3a shows the prefill and decode latency breakdown and speedup from caching under variable context lengths. The speedup increases as cached context gets longer. Furthermore, cache brings more improvements for prefill phase than decode phase, as the fraction in Figure 3b shows. This is because prefill directly benefits from cached context, but decode indirectly benefits from reduced waiting time.
We further analyze the context length in two common LLM tasks where caching can be useful. First, we study the context length of a multi-turn conversation task using a ShareGPT dataset (ShareGPT, 2023; Hugging Face, 2023). Each conversation contains a number of turns between the user and the LLM chatbot. We observe that the context length vary, as Figure 4 shows that 77.2 % of prompts have over 1000 tokens from prior conversation turns. Second, we study another task, document reading comprehension, using the TriviaQA dataset (Joshi et al., 2017). Unlike multi-turn conversation, its context is the whole document that corresponds to the user’s question, with an average context length of 5880 tokens. Therefore, the majority of requests benefit from context caching.
Takeaway 1 Longer context lengths yield more benefits from caching, as more redundant prefill computation can be eliminated. However, context length vary in tasks and prompts, leading to variable benefits. Therefore, longer contexts should be prioritized in caching.
@@ 3.1.2. Impact of Request Rate
The request rate is a key factor that impacts the performance benefits from caching. We conduct an experiment to evaluate the latency under different request rates, as Figure 5 shows. As the request rate becomes higher, the average latency reduction of prefill increases. At the same time, the average decoding latency also reduces and benefits more when the request rate is higher. This is because reduced prefill latency also savings the waiting time for decoding, as discussed in Section 2.2.
@@ 3.1.3. Impact of Cache Size
The cache size is another key factor to performance benefits. We set a fixed request rate of 1.5 prompt/s and vary the cache sizes from 1 to 16 TB. Then, we compare the performance with the no-cache configuration. Figure 6 shows that larger cache sizes yield higher speedup and lower latency, as more prompts hit and reuse the KV cache of the cached context.
Takeaway 3 Larger cache sizes improve performance by increasing hit rates, allowing more prompts to reuse KV caches. However, the benefit may not always scale linearly, as the hit rate improvement slows down once the cache at a certain size.
3.2. Carbon Emissions for LLM Caching
@@ 3.2.1. Cache Carbon Modeling
Section 2.3 introduces the standard carbon modeling method for computing systems. Building upon it, we introduce a carbon modeling technique for LLM caching. For embodied carbon emissions, as caches for LLM context are typically stored in fast storage devices, i.e., SSDs in our study, we model the embodied carbon of cache on top of the storage embodied carbon — it depends on the fraction of storage size used by the cache over the total storage size
(4) $C_{\rm e,Cache}=\frac{\rm S_{util}}{\rm S_{total}}\times\frac{\texttt{T}}{\texttt{LT}}C_{\rm e,SSD}~{},$
where $\rm S_{total}$ is the total size of the SSD, and $\rm S_{util}$ is the size utilized by caching. The same carbon modeling works for other caching mediums such as DRAM, CXL-attached memory, and HDD. We choose SSD because it is a common choice for caching in LLM serving systems (Gao et al., 2025; LMCache Team, 2025b; Yao et al., 2025; Gao et al., 2024; Liu et al., 2024a).
For operational carbon emissions, we first measure the power of server components when running LLM services and the prompt execution latency to obtain energy consumption. Then, we multiply it by the carbon intensity of the grid, as introduced in Equation 2.
Together, the total carbon emission can be calculated using Equation 1. In short, the total carbon is expressed as
(5) $C=E\times\texttt{CI}+\frac{\texttt{T}}{\texttt{LT}}\left(C_{\rm e,Cache}+C_{\rm e,Others}\right)~{}.$
@@
Figure 7. Request rate vs carbon in the ES grid.
We first measure the carbon emissions of the caching system under different request rates. Figure 7 demonstrates that the average per-prompt carbon emission varies with the request rate. We use the carbon intensity of the ES grid to calculate the operational carbon. Cache technique leads to more carbon emissions when request rate is low due to high SSD embodied carbon. As the request rate increases, the operational carbon savings from caching become more prominent, while the relative impact of the SSD’s embodied carbon diminishes.
Takeaway 4 The carbon emission savings depend on request rate. Caching saves more carbon emissions under higher request rates, as caching effectively reduces the high operational carbon under higher loads.
Figure 8. Carbon emission savings from caching in 25 grids. A ratio $<1$ indicates carbon emission reduction.
We next study the impact of carbon intensity. Under a request rate of 1.5 prompt/s, we calculate total carbon emission savings using carbon intensity (CI) from 25 grids (Electricity Maps, 2025), as shown in Figure 8. Grids with a higher CI (on the right side) benefit more from caching. For example, the 16 TB cache reduces 7.5 % carbon emissions in MISO with a high-CI of 485 gCO2e/kWh. However, low-CI grids (on the left side) benefit less from caching and can even lead to higher carbon emissions. For example, the same cache leads to 16.5 % more carbon emissions in FR with a low-CI of 33 gCO2e/kWh. This shift is because the operational carbon dominates the total carbon emission in high-CI grids, while the embodied carbon dominates in low-CI grids. Consequently, when CI is high, a larger cache can save more computations and significantly reduce operational carbon, reducing the total carbon per prompt. When CI is low, embodied carbon takes a larger proportion of the total carbon emissions. Thus, increasing the cache size may negatively impact the carbon emissions.
@@
Figure 9. Carbon emissions per request under different cache sizes and grids.
@@
To better understand this tradeoff between embodied carbon and operational carbon, we examine four grids in Figure 2 and study their per-prompt carbon emission savings under cache sizes ranging from 1 to 16 TB, as shown in Figure 9. In all regions, the fraction of embodied carbon increases as cache size increases, while operational carbon reduction is more significant in high-CI grids. Further, even within the same region, the CI of its grid can change due to renewable energy sources. For example, on May 12, 2025, the CI of CISO can change from 311 at night to 62 at noon (according to Electricity Maps (Electricity Maps, 2025)). As a result, a fixed cache configuration is not the optimal choice under different CIs.
@@
Takeaway 5 Carbon emissions depend on carbon intensities and cache sizes. A higher carbon intensity leads to lower carbon emissions with a larger cache, as the cache saves operational carbon due to computation reduction. When the carbon intensity is low, a large cache can increase the carbon emission due to its embodied carbon.
4. High-level Ideas of EmbAdvisor
@@
In this work, we design a framework EmbAdvisor that enables carbon-aware caching for LLM serving by making a trade-off between the operational carbon reduction from caching and its extra embodied carbon. Next, we introduce two main ideas that enable adaptive caching and guarantee SLO attainment for sustainable LLM serving.
@@ 4.1. Carbon-aware Adaptive Caching
Sections 3.2.2 and 3.2.2 reveal that the carbon savings from caching depend on both request rate and carbon intensity. These two factors affect the choice of cache size configuration dynamically, as they depend on the task, and time and location of the grid. Therefore, the first challenge is to find the optimal cache size given the dynamic factors.
We design an adaptive caching system, as a static configuration cannot handle the dynamics. The high-level idea is to first profile the LLM serving system’s performance under different combinations of request rates and carbon intensity levels. When serving prompts, EmbAdvisor can find the optimal choice of cache size that leads to the lowest per-prompt carbon emissions, according to the current request rate. Figure 10a illustrates an example where EmbAdvisor adaptively increases the cache size when the load (i.e., request rate) increases, to better reduce the operational carbon. Figure 10b shows another example where EmbAdvisor adaptively shrinks the cache size when the carbon intensity reduces, to save the embodied carbon of cache storage.
@@ 4.2. Performance SLO Attainment
Only optimizing for carbon emissions may also negatively impact the performance. For example, a small cache leads to the minimum carbon emissions under low carbon intensity, but significantly degrades performance, as Section 3.1.3 shows. In LLM serving systems, Service Level Objectives (SLOs) typically define latency thresholds for TTFT and TPOT, that needs to be met by the majority of requests (e.g., 90 %) (Zhong et al., 2024; Stojkovic et al., 2025; Qin et al., 2024). Cache configurations that fail to meet these performance requirements are not acceptable, even if they minimize carbon emissions. Therefore, the second challenge to attain SLO while minimizing carbon emissions.
Our approach is to use a constraint solver to address this challenge. We formulate the carbon emission optimization as an Integer Linear Programming (ILP) problem. The solver has an objective function that minimizes the carbon emissions based on the profile discussed in the first high-level idea. At the same time, it enforces constraints that ensure the chosen configuration satisfies the TTFT and TPOT thresholds. Back to the example where a small cache is preferred to minimize carbon but can miss the SLOs, with the constraint solver, EmbAdvisor would choose a larger cache for better performance which attains the SLOs.
5. EmbAdvisor Framework
Figure 11. System overview of EmbAdvisor.
@@ 5.1. System Overview
Based on the aforementioned high-level ideas, we design and implement EmbAdvisor, a carbon-aware caching framework for LLM serving. Figure 11 presents the system overview of EmbAdvisor. EmbAdvisor consists of four major components. First, a profiler periodically analyzes the LLM task to model the relationship between cache size, load, and performance. Second, a performance monitor tracks the latency of LLM serving and the incoming request rate. Third, a constraint solver takes the SLO, carbon intensity, and the current load and performance from the monitor as input, and identifies the cache configuration that minimizes carbon emissions while attaining SLO. Finally, a cache manager that resizes the cache according to the cache configuration. It also incorporates a carbon-saving-aware (CSA) replacement policy that replaces cache entries that lead to the least carbon emissions reductions. In this work, we build our caching system on top of LMCache (LMCache Team, 2025b), which works as the backend of LLM serving.
@@ 5.2. Cache Performance Profiler
The cache performance profiler analyzes the performance and carbon emissions under different request rates and cache sizes for each LLM task. The profiler samples a number of prompts and evaluates them on an initialized cache that has cached prompts reaching its maximum capacity. The profiler consists of two components for evaluation, a monitoring tool and a carbon emission calculator, to generate a profile of each LLM task.
@@
Monitoring. The profiler sweeps cache sizes and request rates. Cache size values are defined by the cache configurations, while request rates are varied up to the maximum level the system can support before violating SLOs. For each combination of cache size and request rate, it tracks the TTFT and TPOT of every prompt. In parallel, the profiler records the power usage of key server components. The CPU power is measured using a RAPL tool (Borghorst, 2018) and GPU power is measured using pyNVML (pyNVML Developers, 2025) every 1 ms. For other relatively low-power components, like SSD and DRAM, that do not provide direct measurement interfaces, the monitoring tool follows the typical power in their specifications (Samsung, 2023; Micron, 2025). In real-world LLM serving scenarios, profiling is performed periodically to adapt to dynamic changes in workload characteristics, such as distribution of request length and access patterns, as seen in prior work (Li et al., 2023c).
@@
Carbon Calculation. The carbon calculator follows the approach in Sections 2.3 and 3.2.1. It first calculates the total energy consumption using the timing and latency measurements and then calculates the operational carbon given the current carbon intensity of the grid, following Equation 2. Next, it calculates the embodied carbon of the platform using Equation 3. Specifically, it calculates the embodied carbon of the SSD for caching using Equation 4. Finally, it calculates the total carbon emissions by summing up the operational and embodied carbon numbers, following Equation 5.
@@
In this work, we adapt two LLM tasks to EmbAdvisor, multiturn-conversation and document reading comprehension, which have also been studied in prior works (Liu et al., 2024a; Zhong et al., 2024; Gao et al., 2024; Liu et al., 2024b; Chen et al., 2025; Luo et al., 2024; Du et al., 2024; Jeong and Ahn, 2025). For each task, we sample 500 prompts from the dataset. We execute these prompts after cache warmup and initialization. We demonstrate the profiling results of both tasks (ShareGPT dataset (ShareGPT, 2023) for multi-turn conversation and TriviaQA (Joshi et al., 2017) for document reading comprehension) using a Llama-3 70B model (Meta, 2024) and platform listed in Table 1. As the profile targets EmbAdvisor, the cache uses our carbon-aware policy (details in Section 5.5). More experiment methodology details are described in Section 6.1.
Figure 12. Profiling results of multi-turn conversation, using a ShareGPT dataset (Hugging Face, 2023) and Llama-3 70B model.
Figure 13. Profiling results of document reading comprehension, using TriviaQA dataset (Joshi et al., 2017) and Llama-3 70B model.
@@
Figure 12 shows heatmaps of the average TTFT and TPOT, and carbon savings from caching, under variable request rates (x-axis) and cache sizes (y-axis), for the multi-turn task. Specifically, the carbon emissions savings are illustrated as the ratio between the no-cache and cached configurations. Thus, a ratio greater than 1 indicates benefits. The TTFT and TPOT in the profile are consistent with Sections 3.1.2 and 3.1.3. Overall, larger caches improve the latency and the benefit is more prominent for higher request rates. The trend of carbon savings is also aligned with Section 3.2.2. Larger caches may not bring carbon emissions savings, especially when the request rate is low. When the rate is high, larger caches yield more savings.
Figure 13 shows the heatmaps of the document reading comprehension task. The observations are aligned with the multi-turn conversation except higher TTFT caused by longer document context as input. Thus, the maximum request rate is lower to meet SLOs. For the same reason, all cases show carbon emission savings, but still, higher request rates benefit more from larger caches. These profiles will be taken as input by a constraint solver (Section 5.4) to select the optimal cache size configuration.
5.3. Performance Monitor
The performance monitor collects statistics from the LLM task, both its load and performance. The load is measured as request rate, i.e., prompt/s. The performance is measured as the TTFT and TPOT, the two critical latencies required by the SLO, in the same way as the cache performance profiler (Section 5.2). The performance monitor sends the statistics to the constraint solver to determine the carbon-optimal cache configuration.

### v2

@@ 3. Performance and Carbon Emission of LLM Caching
In this section, we analyze the performance and carbon emissions of caching in LLM serving and summarize the observations as takeaways. We illustrate the results using the Llama 3 70B model (Meta, 2024) and the ShareGPT dataset (Hugging Face, 2023), evaluated on the server described in Table 1. We use LMCache (LMCache Team, 2025b) as the caching system, which saves context on 16 TB SSDs. The cache has been initialized with 200k prompts, and we record the performance and carbon emissions of 500 prompts. Details of our evaluation methodology are provided in Section 6.1.
3.1. Performance of LLM Caching
We first study the performance of context caching in LLM serving by focusing on two factors: context length and request rate. Since the prefill phase accounts for only a small fraction of the total latency, we present all latency results in log scale to better visualize the prefill latency.
@@ 3.1.1. Impact of Context Length
Caching reduces the computation of the existing context. Therefore, we first evaluate its benefits under different context lengths. Figure 4a shows the prefill and decode latency breakdown and speedup from caching under variable context lengths. The speedup increases as the cached context gets longer. Furthermore, caching brings more improvements for the prefill phase than the decode phase, as the fraction in Figure 4b shows. This is because prefill directly benefits from cached context, but decode indirectly benefits from reduced waiting time.
We further analyze the context length in two common LLM tasks where caching can be useful. First, we study the context length of a multi-turn conversation task using a ShareGPT dataset (ShareGPT, 2023; Hugging Face, 2023). Each conversation contains a number of turns between the user and the LLM chatbot. We observe that the context length varies, as Figure 4 shows that 77.2 % of prompts have over 1000 tokens from prior conversation turns. Second, we study another task, document reading comprehension, using the TriviaQA dataset (Joshi et al., 2017). Unlike multi-turn conversation, its context is the whole document that corresponds to the user’s question, with an average context length of 5880 tokens. Therefore, the majority of requests benefit from context caching.
Takeaway 1 Longer context lengths yield more benefits from caching, as more redundant prefill computation can be eliminated. However, context length varies in tasks and prompts, leading to variable benefits. Therefore, longer contexts should be prioritized in caching.
@@ 3.1.2. Impact of Request Rate
The request rate is a key factor that impacts the performance benefits from caching. We conduct an experiment to evaluate the latency under different request rates, as Figure 6 shows. As the request rate becomes higher, the average latency reduction of prefill increases. At the same time, the average decoding latency also reduces and benefits more when the request rate is higher. This is because reduced prefill latency also saves waiting time for decoding, as discussed in Section 2.2.
@@
Figure 5. (a) Latency of prefill and decode under different request rates, and the speedup from caching. (b) Fraction of prefill and decode latency.
@@ 3.1.3. Impact of Cache Size
The cache size is another key factor for performance benefits. We set a fixed request rate of 1.5 prompts/s and vary the cache sizes from 1 to 16 TB. Then, we compare the performance with the no-cache configuration. Figure 6 shows that larger cache sizes yield higher speedup and lower latency, as more prompts hit and reuse the KV cache of the cached context.
Takeaway 3 Larger cache sizes improve performance by increasing hit rates, allowing more prompts to reuse KV caches. However, the benefit may not always scale linearly, as the hit rate improvement slows down once the cache reaches a certain size.
3.2. Carbon Emissions of LLM Caching
@@ 3.2.1. Cache Carbon Modeling
Section 2.3 introduces the standard carbon modeling method for computing systems. Building upon it, we introduce a carbon modeling approach for LLM caching. In this work, we use SSD, the most commonly used storage in LLM serving systems (Gao et al., 2025; LMCache Team, 2025b; Yao et al., 2025; Gao et al., 2024; Liu et al., 2024b), as the cache hardware. [^1] The same carbon modeling also applies to other caching mediums, such as DRAM, CXL-attached memory, and HDD. In a cloud environment, storage is provisioned on demand and can be flexibly resized. In LLM context caching, the cache size reflects the number of cached tokens, as the service acquires more storage when the number of cached tokens increases and releases storage when the demand for cached tokens decreases. We model the cache’s embodied carbon as proportional to the current allocated storage size $\rm S_{Alloc}$, reflecting a cloud scenario where only the reserved storage contributes to embodied carbon, rather than the entire device. Let $C_{\rm e,SSD}^{\rm Unit}$ denote the embodied carbon per unit of storage (e.g., kgCO2e per TB). The embodied carbon attributed to the allocated SSD is then scaled by the fraction of its lifetime during which the storage is occupied. In our evaluation system, we limit the SSD size to 16 TB, as shown in Table 1. Accordingly, we define the carbon emissions of the cache as follows:
(4) $C_{\rm e,Cache}={\rm S_{Alloc}}\times\frac{\texttt{T}}{\texttt{LT}}C_{\rm e,SSD}^{\rm Unit}.$
For operational carbon emissions, we first measure the power of every server component when serving LLM and the prompt execution latency to obtain energy consumption. Then, we multiply it by the carbon intensity of the grid, as introduced in Equation 2. Together, the total carbon emissions can be calculated using Equation 1, expressed as:
(5) $C=E\times\texttt{CI}+\frac{\texttt{T}}{\texttt{LT}}\left(C_{\rm e,Cache}+C_{\rm e,Others}\right).$
@@
Figure 7. Carbon emissions per request (a) in the ES grid under different request rates, and (b) under different cache sizes and grids (based on grid average CI in Figure 2a).
We first measure the caching system’s carbon emissions under varying request rates. Figure 7(a) demonstrates that the average per-prompt carbon emissions vary with the request rate. All carbon emissions are calculated with ES grid’s carbon intensity. At low request rates, caching incurs higher emissions due to the embodied carbon of SSDs. As the request rate increases, operational carbon savings from caching become dominant, while the relative impact of embodied carbon diminishes.
Takeaway 4 The carbon emission savings depend on the request rate. At higher loads, caching yields greater carbon reductions by effectively mitigating the elevated operational carbon.
Figure 8. Carbon emission savings from caching in 12 grids. A ratio ¡ 1 indicates carbon emission reduction.
We next study the impact of carbon intensity. Under a request rate of 1.5 prompts/s, we calculate the total carbon emission savings using the average carbon intensity (CI) from 12 grids (Electricity Maps, 2025), as shown in Figure 8a. Grids with a higher CI (on the right side) benefit more from caching. For instance, a 16 TB cache reduces carbon emissions by 7.5 % in MISO, where the CI reaches 485 gCO2e/kWh. In contrast, low-CI grids (on the left side) benefit less or even incur higher emissions from caching. For example, in FR, where the CI is only 33 gCO2e/kWh, the same cache increases carbon emissions by 16.5 %. As CI fluctuates within the same grid due to changing energy sources, we next evaluate the carbon emission savings of a 16 TB cache over a day in the CISO grid, as shown in Figure 8b. The CI drops to its daily minimum of 37 gCO2e/kWh at 7 AM when the grid is mainly powered by renewable energy sources (as shown in Figure 2b), making caching not carbon-efficient. Conversely, at 8 PM, CI peaks at 232 gCO2e/kWh, and caching achieves its maximum carbon reduction.
The shift in carbon emission savings results from changes in the primary carbon contributor within the LLM serving system. Consequently, when CI is high, a larger cache can save more computations and significantly reduce operational carbon, reducing the total carbon per prompt. Conversely, at low CI, embodied carbon dominates the total carbon emissions. Thus, increasing the cache size may negatively impact the carbon emissions.
@@
To better understand this tradeoff between embodied carbon and operational carbon, we further evaluate the carbon emission savings under cache sizes ranging from 1 to 16 TB with the average CI of four grids in Figure 2, as shown in Figure 7(b). In all grids, the fraction of embodied carbon rises with cache size, while operational carbon reduction is more significant in high-CI grids. These results indicate that a fixed cache configuration is suboptimal under variable CIs.
@@
Takeaway 5 Carbon emissions depend on carbon intensities and cache sizes. A higher carbon intensity can lead to lower carbon emissions with a larger cache, as the cache saves operational carbon due to computation reduction. When the carbon intensity is low, a large cache can increase the carbon emissions due to its embodied carbon. Because the carbon intensity is highly dynamic, both across grids and within a grid, an adaptive method is needed to achieve optimal carbon savings.
4. High-level Ideas of GreenCache
@@
GreenCache is a carbon-aware caching framework for LLM serving by making tradeoffs between the operational carbon reduction from caching and its extra embodied carbon. Next, we introduce two main ideas that enable adaptive caching and guarantee SLO attainment behind it.
@@ 4.1. Carbon-aware Adaptive Caching
Sections 3.2.2 and 3.2.2 reveal that the carbon savings from caching depend on both request rate and carbon intensity. These two factors affect the choice of cache size configuration dynamically, as they depend on the task, time, and the grid. Therefore, the first challenge is to find the optimal cache size given the dynamic factors.
GreenCache resizes the cache according to the request rate and carbon intensity during runtime. The high-level idea is to first profile the LLM serving system’s performance and power under different combinations of request rates and carbon intensity levels. When serving prompts, GreenCache can find the optimal configuration of cache size that leads to the lowest per-prompt carbon emissions. However, one challenge is that cache resizing, especially enlarging the cache, takes time to warm up, and thus the performance highly depends on the future conditions. To overcome this challenge, GreenCache predicts the request rate and carbon intensity based on historical data, enabling optimal decision-making that considers not only the current condition but also future trends. Figure 10a illustrates an example where GreenCache adaptively increases the cache size when the predicted load (i.e., request rate) increases to better reduce the operational carbon. Figure 10b shows another example where GreenCache adaptively shrinks the cache size when the carbon intensity is predicted to reduce, to save the embodied carbon of cache storage.
@@
Figure 10. System overview of GreenCache. Components in green are newly designed by this work.
@@ 4.2. Performance SLO Attainment
Only optimizing for carbon emissions can also negatively impact performance. As Section 3.1.3 has shown, a small cache can lead to the minimum carbon emissions under low carbon intensity, but significantly degrade performance. In LLM serving systems, Service Level Objectives (SLOs) typically define latency constraints for TTFT and TPOT, with the goal of ensuring that a high percentage of requests (e.g., 90 %) adhere to these thresholds. Cache configurations that fail to achieve the required SLO attainment are not acceptable, even if they minimize carbon emissions. Thus, the second challenge lies in minimizing carbon emissions while ensuring SLO attainment.
GreenCache uses a constraint solver to find such a cache configuration. We formulate the constrained carbon emission optimization as an Integer Linear Programming (ILP) problem. The solver has an objective function that minimizes the carbon emissions based on the profile discussed in the first high-level idea. At the same time, it enforces constraints to ensure SLO attainment. In scenarios where a small cache is preferred to minimize carbon but can miss most SLOs, the solver guides GreenCache to choose a larger cache that achieves targeted SLO compliance.
5. GreenCache Framework
We first present an overview of GreenCache and then describe its components.
@@ 5.1. System Overview
We design and implement GreenCache, a carbon-aware caching framework for LLM serving. Figure 10 presents the system overview of GreenCache. GreenCache consists of six major components. First, a profiler periodically analyzes the LLM task to model the relationship between cache size, load, performance, and power consumption (details in Section 5.2). Second, a performance monitor tracks the TTFT and TPOT of LLM serving and the current load (i.e., request rate). Third, a combination of CI predictor and load predictor forecasts the CI and request rate based on historical values. Specifically, we use the state-of-the-art CI predictor, EnsembleCI, (Yan et al., 2025) for CI prediction (methodology details in Section 6.1), and design a load predictor (details in Section 5.3). A constraint solver takes the SLO, predicted CI, performance, and power profile, predicted load, and current performance as input, and identifies the cache configuration that minimizes carbon emissions while attaining SLO (details in Section 5.4). Finally, a cache manager resizes the cache according to the cache configuration. It also incorporates a carbon-aware replacement policy that replaces cache entries with the Least Carbon Savings (LCS) (details in Section 5.5). We build GreenCache on top of the LLM caching system, LMCache (LMCache Team, 2025b). GreenCache can be flexibly adapted to different LLMs.
We next introduce the new components in the GreenCache framework, marked green in Figure 10.
@@ 5.2. Cache Performance Profiler
The cache performance profiler analyzes performance and power under different request rates and cache sizes for each LLM task. The profiler samples a number of prompts and evaluates them on an initialized cache filled to maximum capacity. The profiler consists of two components for evaluation, a monitoring tool and a carbon emission calculator, to profile each LLM task.
@@
Performance and Power Analysis. The profiler sweeps cache sizes and request rates. Cache size values are defined by the cache configurations, while request rates are varied up to the maximum level the system can support before violating SLOs. For each combination of cache size and request rate, it records the TTFT and TPOT of every prompt. In parallel, it records the power consumption of key server components. The CPU power is measured using a RAPL tool (Borghorst, 2018) and GPU power is measured using pyNVML (pyNVML Developers, 2025) every 1 ms. For other relatively low-power components, like SSD and DRAM, that do not provide direct measurement interfaces, the monitoring tool follows the typical power in their specifications (Samsung, 2023; Micron, 2025). In real-world LLM serving scenarios, profiling is performed periodically to adapt to dynamic changes in workload characteristics, such as distribution of request length and access patterns, as seen in prior work (Li et al., 2023c).
@@
Carbon Calculation. The carbon calculator follows the approach in Sections 2.3 and 3.2.1. It first computes the total energy consumption using the timing and latency measurements and then derives the operational carbon based on the grid’s current carbon intensity, as defined in Equation 2. Next, it estimates the embodied carbon of the platform using Equation 3. Specifically, it calculates the embodied carbon of the SSD for caching via Equation 4. Finally, the total carbon emissions are obtained by summing the operational and embodied components, following Equation 5.
Figure 11. Profiling results of TTFT and TPOT (lower is better), and carbon savings over no-cache (higher is better) when serving Llama-3 70B in the ES grid for (a) multi-turn conversation, using a ShareGPT dataset (Hugging Face, 2023) and (b) document comprehension using a TriviaQA dataset (Joshi et al., 2017) (with skewness of $\alpha=0.4$).
@@
In this work, we adapt two LLM tasks to GreenCache, multi-turn conversation and document reading comprehension, which have also been studied in prior works (Liu et al., 2024b; Zhong et al., 2024; Gao et al., 2024; Liu et al., 2024a; Chen et al., 2025; Luo et al., 2024; Du et al., 2024; Jeong and Ahn, 2025). For each task, we sample 500 prompts from the dataset and execute them after cache warmup. We demonstrate the profiling results of both tasks (ShareGPT dataset (ShareGPT, 2023) for multi-turn conversation and TriviaQA (Joshi et al., 2017) for document reading comprehension) using a Llama-3 70B model (Meta, 2024) on the platform described in Table 1. As the profile targets GreenCache, the cache uses our carbon-aware policy (details in Section 5.5). More experiment methodology details are described in Section 6.1.
@@
Figure 11(a) shows heatmaps of the average TTFT and TPOT, and carbon savings from caching, under variable request rates (x-axis) and cache sizes (y-axis), for the multi-turn conversation task. We profile TTFT and TPOT under various cache sizes (y-axis). We also profile the carbon emission savings based on the CI of the ES grid — defined as the ratio between the no-cache and cached configurations — where a ratio greater than 1 indicates benefits. The TTFT and TPOT in the profile are consistent with Sections 3.1.2 and 3.1.3. Overall, larger caches improve the latency, and the benefit is more prominent for higher request rates. The trend of carbon savings is also aligned with Section 3.2.2. Larger caches may not bring carbon emission savings, especially when the request rate is low. When the rate is high, larger caches yield more savings. Figure 11(b) shows the heatmaps of the document reading comprehension task with skewness of $\alpha=0.4$ (i.e., 10 % of documents are accessed by ~25 % of prompts). The latency and carbon savings trends are consistent with the multi-turn conversation task. However, the longer document contexts result in a TTFT, which lowers the maximum request rate in this profile. These profiles will be used by a constraint solver (Section 5.4) to select the optimal cache size configuration.
5.3. Load Predictor
Prior work from Microsoft has shown that LLM prompts follow a similar pattern in a day (Stojkovic et al., 2025), where the rate mostly depends on the time of day. We design a lightweight predictor, based on the Seasonal Autoregressive Integrated Moving Average (SARIMA) model, using pmdarima (Smith and others, 2017). It captures both daily periodicity and short-term autocorrelation in the given load. The SARIMA parameters are estimated from historical rate traces using maximum likelihood fitting. We also use pmdarima for parameter auto-tuning. We use a hold-out evaluation, where it takes the most recent three consecutive days of data as input and predicts 24 hours ahead. Moreover, the load predictor performs online step-ahead prediction during runtime — every hour, the model incorporates the most recent load before forecasting future hours to adapt to rate fluctuations. This predictor incurs minimum computation and runs on separate CPU cores, without interfering with the LLM workload.